OpenStudy (anonymous):

change into y=mx+b formula 2x+3y<6 5x-3y<15

OpenStudy (anonymous):

2x + 3y < 6 3y < 6 - 2x OR 3y < -2x + 6 y < -2/3X + 6/3 y < -2/3X + 2 \[y = -\frac{ 2 }{ 3 } X + 2\] 5x - 3y <15 -3y < 15 - 5x OR -3y < -5x +15 y > 5/3 - 15/3 - Notice Inequality signs change when you divide/multiply a negative number across them \[y > \frac{ 5 }{ 3 } X - 5\]

OpenStudy (anonymous):

thank you! where do i plug in (0,0) for a test point to graph?

OpenStudy (anonymous):

I don't understand what you are asking.

OpenStudy (anonymous):

the original equations? 2x+3y<6 & 5x-3y<15

OpenStudy (anonymous):

What about them?

OpenStudy (anonymous):

i have to graph the equation and shade. so my test point is (0,0)

OpenStudy (anonymous):

does that get plugged into my original equations?

OpenStudy (anonymous):

Do you know how to graph them?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Why would you plug in the point (0,0) ?

OpenStudy (anonymous):

(0,0) is called the origin. It's where The X-axis and Y-axis intersect. Neither one of these lines pass through the origin.

OpenStudy (anonymous):

the two systems are linear inequalities. since the two lines don't pass through (0,0) you can use that as a test point to check that your equation is true. i just found it in my notes. so you plug in 2(0)+3(0)<6 0<6 is true. then you do 5(0)-3(0)<15 0<15 is also true so you shade everything that satisfies the equation through your test point.

OpenStudy (anonymous):

Ah ok, nobody ever asks me to shade anymore, but I briefly remember that from many years ago.